Write the equation of the line passing through (6,4) and (6,2).
Answer by checkley71(8403)

SLOPE=(Y2-Y1)/(X2-X1)=(2-4)/(6-6)=-2/0 WHICH INDICATES A VERTICLE LINE THAT PASSES THROUGH THE POINT
THEREFORE THE LINE EQUATION IS X=6.